OpenStudy (anonymous):

Suppose that the augmented matrix of a system of linear equations for unknowns x, y, and z is [1 0 -3 | -(9/4) [4 1 -7 | -13 [-3 0 -9 | (27/4) Find the solution. Please help ASAP. P.S- Sorry about the formatting. I could not figure out how to copy/paste the problem so I had to try and make it work haha. THANK YOU!!!

OpenStudy (anonymous):

Also, I made a typo. In the last row it should be (9) and not (-9)

OpenStudy (anonymous):

SOLVE THE SYSTEM. PLEASE HELP

OpenStudy (sidsiddhartha):

reduce it to echelon form

OpenStudy (sidsiddhartha):

u know what is echelon matrix?

OpenStudy (anonymous):

I know that I have to do that but I am not sure how to. I need this answer quickly! Can you tell me if this is correct or not? x=-(9/4) y=-4 z=0

OpenStudy (sidsiddhartha):

ya the echelon matrix is 1 0 -3 -9/4 0 1 5 -4 0 0 0 0 if u solve it z=0, y=-4,x=-9/4

OpenStudy (anonymous):

Are you 100% positive? I have to get this correct and I just want to make sure. Also, I just want to make sure that you saw how i said the third number in the last row is 9, not -9. I just want to be sure haha

OpenStudy (sidsiddhartha):

i think its ok

OpenStudy (jack1):

attached if it helps...?

OpenStudy (anonymous):

That helps a lot! So when I write my answer would I write x = -9/4 + 3z y = -4 - 5z z = 0 or just x = -9/4 y = -4 z = 0 ??

OpenStudy (jack1):

depends if u consider z to be zero or unknown, if z = 0 then both are correct, if z = unknown, then only the top series of answers is correct

OpenStudy (anonymous):

Okay! Thanks so much everybody!